an isa 200×100×10 angle is welded to steel plate by means of fillet welds the angle is subjected to static force of 150 kN and the permissible shear stress for the weld is70kN. determine length of welds at top and bottom .l1=100mm,and l2=200mm
1
Expert's answer
2020-06-05T05:36:17-0400

Given static loading =150 kN = 150 "\times 10^3" N ,here dimension is given as 200 "\times 100\times 10" ,

permissible shear stress= 70 "\frac{kN}{mm^2}" ,here a+b=200


s= size of weld = thickness= 10 ,we know that

for single weld

"150 \times 10 ^3=0.707 \times s \times l\times \tau"

on putting value we can get value of l

l= 303.09 mm

"l_a+ l_b=303.09"

now distance of centroidal axis from bottom angle

b=55.3 and a= 144.7 so

"l_a= \frac{l\times b}{a+b}"

"l_a=104.2" and "l_b= 198.89"